Mary's demand curve for food is Q =10


Her price elasticity of demand for food at price p* equals - 2/3 how much is p*



Note


a Marginal Cost (MC) is the derivative of the total cost (TC) with respect to quantity Q. TC = 300, + 5Q +10Q2 Then MC= 5 + 20Q



1
Expert's answer
2021-12-08T19:24:21-0500

Solution:

Price elasticity of demand = % change in quantity demanded /% change in price

PED = "\frac{\triangle Q}{\triangle P} \times \frac{\triangle Q}{\triangle P}"


"\frac{\triangle Q}{\triangle P}" = -0.25

Q = 10


"-\frac{2}{3} =-0.25\times \frac{P}{10}"


P = 26.67

Price p* equals = 26.67